The following data pertain to products A and B, both of which are purchased by Madame X.

Initially, the prices of the products and quantities consumed are:

PA = $8, QA = 5, PB = $6, QB = 10

Madame X has $100 to spend per time period. After a reduction in the price of B, the prices and quantities consumed are:

PA = $8, QA = 4, PB = $9, QB = 7.5

Assume that Madame X maximizes utility under both price conditions above. Also, note that if after the price reduction enough income were taken away from Madame X to put her back on the original indifference curve, she would consume this combination of A and B:

QA = 10, QB = 4

1) The change in consumption of good B due to the substitution effect is ____________ and the change in consumption of good B due to the income effect is________ .

2) Determine if the product B is a normal, inferior, or Giffen good
Answer #1

The total effect of price change is the difference in quantities before and after the price change is 7.5-10=-2.5 which includes both income and substitution effect. The reduction in consumption because of reduction in income is income effect which is 7.5-4=3.5
So, the difference between -2.5-3.5 = -6, represents the substitution effect

1) The change in consumption of good B due to substitution effect is -6 and the change in consumption of good B due to the income effect is 3.5
2) The product is an normal good as the with the decrease in income the consumption decreases and increase in price leads to decrease in consumption